OpenStudy (anonymous):

HELP PLEASE :) Two concentric spheres have radii of 5" and 6". Find the volume of the space between them. cubic inches

OpenStudy (anonymous):

Now tell me the formula for Volume

OpenStudy (anonymous):

The Rest is Easy

OpenStudy (anonymous):

i know i cant fgure out what to do though

OpenStudy (anonymous):

it's okay let me give you some hints

OpenStudy (anonymous):

\[Volume_{Sphere} = \frac{4}{3}\pi r^3 \]

OpenStudy (anonymous):

So now just put in the Value Now the Remaining Volume \[Re_{vol} = \frac{4}{3}\pi(6^3 - 5^3)\]

OpenStudy (anonymous):

That is your answer